2. Suppose you deposit $100 million in the hank. How many years later you will withlraw $200 million if annual interest rate is 2.5%?
Answer #1

Let the years be n in number

FV = PV x (1 + r)^n

200 = 100 x (1 + 2.5%)^n

2 = 1.025^n

ln(2) = n x ln(1.025) (taking natural logs)

This gives n = ln(2)/ln(1.025) = 29 years minimum

Thus, it takes 29 years for 100 million to become 200 million at 2.5%